40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 126 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  10976   Wed Feb 4 19:21:45 2015 JenneUpdatePEMSeismometer Vertex is covered

I opened up the spaghetti pot over the vertex seismometer, and taped the cable to the slab.  The way the cable is coiled, it was touching the underside of the seismometer.  Now the only connection is at the cable connector.  There is a ~few inch bit of cable, then it's taped down. 

  10975   Wed Feb 4 19:21:37 2015 KojiUpdateASCArm ASS servos now have triggered gain with arm lock status

We had persistent frustration by occasional unlock during ASSing.
Today, I added triggers to the servo gains in order to elliminate this annoyance.

Each ASS servo gain slider is multiplied with the corresponding LSC Trigger EPICS channel (i.e. C1:LSC-iARM_TRIG_MON, where i=X or Y).
This has been done by ezcaread modules in RCG.

The model and screen have been commited to svn.

  10974   Wed Feb 4 18:27:55 2015 KojiSummaryASCXarm ASS fix

Please remember that Xarm ASS needs FM6 (Bounce filters) to be ON in order to work properly.

  10973   Wed Feb 4 18:16:44 2015 KojiUpdateLSCData transfer rate of c1lsc reduced from ~4MB/s to ~3MB/s

c1lsc had 60 full-rate (16kS/s) channels to record. This yielded the LSC to FB connection to handle 4MB/s (mega-byte) data rate.
This was almost at the data rate limit of the CDS and we had frequent halt of the diagnostic systems (i.e. DTT and/or dataviewer)

Jenne and I reviewed DAQ channel list and decided to remove some channels.  We also reviewed the recording rate of them
and reduced the rate of some channels. c1lsc model was rebuilt, re-installed, and restarted. FB was also restarted. These are running as they were.
The data rate is now reduyced to ~3MB nominal.


The following is the list of the channels removed from the DQ channels:

AS11_I_ERR
AS11_Q_ERR
AS165_I_ERR
AS165_Q_ERR
POP55_I_ERR
POP55_Q_ERR

The following is the list of the channels with the new recording rate:

TRX_SQRTINV_OUT 2048
TRY_SQRTINV_OUT 2048
DARM_A_ERR 2048
DARM_B_ERR 2048
MICH_A_ERR 2048
MICH_B_ERR 2048
PRCL_A_ERR 2048
PRCL_B_ERR 2048
CARM_A_ERR 2048
CARM_B_ERR 2048

  10972   Wed Feb 4 14:30:05 2015 ericqUpdateLSCASDC Whitening Gain

At the lunch meeting, we were thinking about the exess high frequency content of the MICH control signal, which leads to railing against the FM output limiter and lock loss. I propsed that POPDC sensor/ADC noise was the culprit. 

In short, I was wrong. Just now, I locked the PRMI with a MICH offset as we normally do, and then froze the POPDC output. The MICH spectrum did not change in any noticible way. 

However, increasing the analog ASDC whitening gain showed a direct improvement of the error signal noise floor, and thus a reduction in the control signal spectrum. I.e. we have been suffereing from ASDC ADC noise.

We had previously set the ASDC whitening gain so that half fringe of the PRMI would be well within the ADC range, but since we're actually only ever going to 25%, I feel fine upping this gain. Interestingly, when increasing the whitening gain by 12dB,  the control signal spectrum has fallen by more like 20 dB in the 400Hz-1kHz region, which is great. 

The only potential hurdle I can think of is that when we start reducing the MICH offset at zero CARM offset, we may approach ADC saturation in ASDC before we can hand off to RF signals, in which case we would have to dynamically lower the whitening gain, which introduces offsets, which could get hairy. But, since MICH railing has been directly seen to lead to lock-loss, I'd rather solve that problem first. 

  10971   Wed Feb 4 04:51:14 2015 diegoUpdateLSCCARM Transition to REFL11 using CM_SLOW Path

[Diego, Jenne, Eric]

Tonight we kept on following our current strategy for locking the PRFPMI:

  • the first few tries were pretty unsuccessful: the PRMI lock wasn't much stable, and we never managed to reduce CARM offset to zero before losing lock;
  • then we did some usual manteinance: we fixed the X arm green beatnote, fixed the phase tracker and given much attention to ASS alignment, since the X arm wasn't doing great;
  • the last few locks were consintently better: we managed to get to CARM offset zero "easily", but the arm power was not very high (~8);
  • then we tried to transition CARM to REFL11, both with the usual configuration and using CM_SLOW, using REFL11 as input for the Common Mode Board;
    • with the usual configuration, we lost lock right after the transition, because of MICH hitting the rail;
    • we did a very smooth CARM transition directly to REFL11 on the CM_SLOW path; we managed to take a spectrum with the SR785, but we couldn't take any more measurements before losing lock because of some weird glitch, as we can see from the lockloss plot;
  • another thing that helped tonight was changing the ELP of the MICH filter bank: it went from 4th order to 6th order, and from 40 dB suppression to 60 dB suppression;

both of the last two locks, the most stable ones (one transition to usual REFL11 and one transition to "CM_SLOW" REFL11) were acquired actuating on MC2;

 


EDITs by JCD:  At least one of the times that we lost PRMI lock (although kept CARM and DARM lock on ALS), was due to MICH hitting the rail, even after we increased the limiter to 15,000 counts.


 Here is the transfer function between CARM ALS (CARM_IN1) and REFL11 coming through the CM board (CARM_B), just before we transitioned over.  Coherence was taken simultaneously as usual, I just printed it to another sheet.

CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS.pdf

CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS_Coh.pdf


Here is the lockloss plot for the very last lockloss.  This is the time that we were sitting on REFL11 coming through the CM_SLOW path.  A DTT transfer function measurement was in progress (you can see the sine wave in the CARM input and output data), but I think we actually lost lock due to whatever this glitch was near the right side of the plots.  This isn't something that I've seen in our lockloss plots before.  I'm not sure if it's coming from REFL11, or the CM board, or something else.  We know that the CM board gives glitches when we are changing gain settings, but that was not happening at this time.


Q: Here's the SR785 TF of CARM locked through CM board, but still only digital control; nothing exciting. Excitation amplitude was only 1mV, which explains the noisy profile. 

Attachment 1: CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS.pdf
CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS.pdf
Attachment 2: CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS_Coh.pdf
CARM_3Feb2015_CarmBwasCMslow_CarmAwasLiveALS_Coh.pdf
Attachment 3: Glitch_in_CARM_and_PRCL_3Feb2015.png
Glitch_in_CARM_and_PRCL_3Feb2015.png
Attachment 4: slowCM_04-02-2015_042805.png
slowCM_04-02-2015_042805.png
  10970   Tue Feb 3 16:43:48 2015 manasaUpdateGeneralRF amplifier for ALS

I pulled out the RF amplifier box from the IOO rack again and added the new RF amplifiers for FOL. 

I replaced one of the decoupling capacitors of the ALS beat note RF amplifier ; the poloproylene capacitor with a ceramic capacitor (0.1uF) .

After putting back the box, I confirmed that we had a beat note. I did not get a chance to measure the ALS noise after putting back the box because the IFO was already occupied.

I will post a detailed elog of the components in the RF amplifier box once I am done with it (hopefully tomorrow)!

  10969   Tue Feb 3 16:36:33 2015 ericqUpdateLSCCM servo & AO path status

I have removed REFLDC and the SR560 offsetter from the CM board IN2. Now, analog AS55 I lives there, for our single arm testing. (Analog I has more of the single arm Y PDH signal in it). REFL11 has been reconnected to IN1. 


With ITMX super misaligned, Diego and I locked the Y-arm with the AO path on AS55, ultimately at 4kHz bandwidth, but with plenty of gain margin. We didn't allocate the gains too intelligently, and had the CM board input gain slider maxed out, but plenty of headroom in the digital and AO sliders, making it inconvenient to up the UGF even more, to engage the super boosts. However, since this is just a test case to make sure we still can AO lock, I'm not too worried about this. 

Since LSC FMs and such had changed around, old recipies didn't neccesarily work 1:1. Diego is writing a script for the current recipe, and will post an elog with the steps. 

Gains and signs are able to be tracked by loop TFs, the real sticking point is a stable crossover. We used the 1.6k:80 hardware filter in the CM board to give the AO Path a 1/f shape in the crossover region, and undid it digitally in the CM_SLOW input FM. However, we do use a 300:80 in the MC2 sus FM to make the digital loop like 1/f^2 around the crossover, once a little bit of AO has come in to pull up the digital loop's phase. We used the CARM filter bank to do this, so I think we should be able to use a similar technique to do it in the PRFPMI case, as long as the coupled cavity pole is around ~100Hz. 

Attached are a few OLTFs from the progression.

Attachment 1: yarmAO.pdf
yarmAO.pdf
  10968   Tue Feb 3 15:20:13 2015 Steve, KojiUpdateVACPneumatic pressure is being read again

[Koji, Steve]

Summary

The N2 pressure reading (C1:VAC-N2PRES) is now up-to-date after rebooting c1vac1.
The vaccum system is "Vacuum normal". We now have a space pressure transducer.

Introduction

Our vacuum valves are manipulated with 60~75 PSI of nitrogen. All the valves are configured to be closed in the case of low N2 supply pressure.
In order to avoid this safety shutdown accidentally triggered, we have two N2 cylinders to sustain the vacuum valves. When one cylinder goes to low
the mechanical valve switches over to the other cylinder.

We have the monitor channel for this (combined) cylinder pressure. One shoulbe be able to see periodical pressure variation when the auto cylinder
switch is operating. However, the nirogen pressure reading got stuck at 66 PSI on Dec.16, 2014 (See attached 60-day plot of N2 supply pressure). 

What we did

This morning we tracked down the cause of the trouble. We first closed the valves on EPICS and started to vary the N2 pressure.

Our first guess was the pressure transducer (Omega #236PC100GW) that was already 15 yrs old. We even has a sensor spare for replacement.
But it turned out that the direct voltage reading (to be 1mV/PSI) is changing correctly. The second guess was Omega Controller-Monitor
#DPiS32-C24 that is reading the voltage from the tranceducer. The display on this small black unit was changing corresponding to the
pressure change.

So our thought was
1) RS232C of the monitor unit is not working correctly
or
2) c1vac1 is not communicating with the monitor unit.

We wondered what could cause c1vac1 not communicating with the monitor unit, but we were afraid that some function got stuck
during either the nodus upgrade or chiara rebooting (or something else). So we decided to reboot c1vac1

In order to avoid any glitch in the main vacuum pressure, Steve disconnected some of the controller connectors for the closed valves.
We did this treatment before and it was successful.

Then c1vac1 was rebooted just by telnet and type reboot in the terminal.
Once the target is back in action, we noticed that the monitor value started to move.

Steve reverted the cables to the valves and operated the valves to recover "Vacuum Normal" state. Everything is now nicely settled.

Attachment 1: N2pressureReadingIsBack.png
N2pressureReadingIsBack.png
  10967   Tue Feb 3 04:07:49 2015 JenneUpdateModern ControlMeasured PRM -> POP QPD TFs

Today (after centering the POP QPD), I measured the PRM to POP QPD transfer functions.  I am suspicious that this was part of my problem last week.  Since most of the angular noise is coming from the folding mirrors, but I can't actuate on them, I need to know (rather, the Wiener calculator needs to know) how actuating on the PRM will affect the spot at the POP QPD.

For the plots below, I have cut out any data points that have coherence less than 0.95. I may want to go back and fill in a little bit some of the areas (particularly around 3Hz) that I had trouble getting coherence in.

Using these to prefilter my witness data, I am failing to calculate a Wiener filter.  I have tried the Levinson algorithm, as well as brute-forcing it, but I'm too close to singular for either to work.  I am able to calculate a set of Wiener filters without the prefiltering, or with a dummy very simple prefilter, so it's not inherently in the calculators.  Separately, I can plot my vectfit-ed actuator TFs, and I can convert them to a discrete fiilter with the bilinear transform, and then use sosfilt to filter some white noise data, which comes out with the shape I expect.  So.  It's not inherently the filters either.  More work to do, when it's not 4am.

Here are the measured actuator transfer functions.  They were measured as usual with DTT, but since the measurement kept getting interrupted (MC unlock, or I wanted to add more integrations or more cycles), these are several different DTT files stitched together.  In both cases I am acuating PRM's ASC[pit, yaw] EXC point, and looking at the POP QPD.

 

Attachment 1: PIT_measured_TF.png
PIT_measured_TF.png
Attachment 2: YAW_measured_TF.png
YAW_measured_TF.png
  10966   Tue Feb 3 04:01:55 2015 diegoUpdateLSCCM servo & AO path status

[Diego, Jenne]

Tonight we worked on the CM board and AO path:

  • at first we changed the REFL1 input to the CM board from REFL11 to AS55, as written in my previous elog; we tried following Koji's procedures from http://nodus.ligo.caltech.edu:8080/40m/9500 but we didn't get any result: we could lock using the regular digital path but no luck at all for the analog path;
  • then we decided to follow the procedure to the letter, using POY11Q as input to the CM board;
    • we still couldn't lock following the Path #2, even after adjusting the gains to match the current configuration for the Yarm filter bank;
    • we had some more success using Path #1, but we had to lower the REFL1 Gain to ~3-4 (from the original 31) because of the different configuration of the Yarm filter bank, in order to have the same sensing in both of them; we managed to acquire lock a few times, it's not super stable but it can keep lock for a while;
    • when we tried to increase the gain of the MC filter bank and the AO Gain, however, we immediately had some gain peaking, and we couldn't go further then 0.15 and 9db respectively. We currently don't have an answer for that.
    • anyhow, we took a few measurements with the SR785:

 

The BLUE plot is at MC Gain = 0.10 and REFL1 Gain = 4dB; the GREEN plot is for MC Gain = 0.10 and REFL1 Gain = 3dB, which seemed a more stable configuration; after this last configuration, we increased the MC Gain to 0.15 and the AO Gain from 8dB to 9dB and took another measurement, the RED plot; this is as far as we got as of now. We also couldn't increase the REFL11 Gain because it made things unstable and more prone to unlock.

So, some little progress on the AO path procedure, but we are very low on our UGF and we have to find a way to increase our gains without breaking the lock and avoiding the gain peaking we have witnessed tonight.

 

Notes:

  • is the REFL1 Gain dB slider supposed to go to negative dBs? During the night we also tried to use negative dBs, but it seemed it wasn't doing anything instead;
  • when we plugged POY11Q to the CM board, we noticed that it wasn't connected to anything at the moment; since we phase rotate POY11, we were assuming that we were using that signal somewhere. We are confused by this...
  • we remind that REFL11 is no more connected to the CM board input, as POY11 is.
Attachment 1: CARM_03-02-2015_031754.pdf
CARM_03-02-2015_031754.pdf
  10965   Mon Feb 2 22:59:49 2015 diegoUpdateLSCCM board input switched to AS55

[Diego, Jenne]

We just changed the input to the CM board from REFL11 to AS55.

 

  10964   Mon Feb 2 16:58:56 2015 manasaUpdateGeneralRF amplifier for ALS

Today I was around the IOO rack.

I shutdown the power to the beat PDs on the PSL table and disconnected the D sub 3w3 connector that was powering the RF amplifier panel on the IOO rack.

I moved the RF amplifier from the panel to a box that can go on the rack. The box will also hold the RF amplifiers that will be used for FOL. I have not completed putting in all the amplifiers. But the RF amplifier for ALS is in place and the box has been installed on the IOO rack for locking tonight. The power supply to the green beat PDs has been switched ON.

I took the out of loop noise measurements for ALS X and Y and the attachment is the screenshot of it (X and Y have rms of ~300Hz and ~400Hz).

I had to touch the Y end steering mirrors for green to get maximum GTRY before making thes measurments.

Attachment 1: ALSXY_outLoop.png
ALSXY_outLoop.png
  10963   Mon Feb 2 12:24:27 2015 JenneUpdateASCPOP QPD centered

After aligning the PRC, I centered the POP QPD.

  10962   Sat Jan 31 01:34:12 2015 JenneUpdateLSCNot able to engage AO path

Nothing earth-shattering today.

A few things of note:

  • I checked the coherence (no lock present, just noise) between REFL11_I_IN1 and CM_SLOW_OUT, which are meant to be the same thing when only the REFL1 path is allowed through the CM board.
    • However, at first, there was very little coherence - small band, and only about 0.7 or so.
    • I went inside and jiggled the cables, and also toggled the whitening switches for REFL11 and the CM_SLOW, and after that I had excellent coherence. 
      • I didn't take a coherence spectrum in between those activities, but since the cable connections were all solid, I believe that it may have been a sticky slider -esque problem, and the CM whitening wasn't matched between the analog and digital.
    • I tried two or three times to engage the AO path, but I always lost lock before I was getting any meaningful gain through.
      • I took some TFs remotely with the SR785, but they're totally noise.  I don't even know which sign of the CM board is correct, so no real knkowledge added there, from today.
  • The ~400Hz ringing that we have been seeing, we have been blaming on the PRCL loop.  However, just before my last lockloss I saw gain peaking around 400Hz in the CARM input spectrum. I didn't see if it was also there in the PRCL spectrum.  So, either it is coupling from PRCL to CARM, or CARM itself.  But I think we need to see if we can eek out a little more phase at higher frequencies for both of those loops.  I  just looked, and about 16 seconds before the last lockloss, I see the PRCL loop coupling into the CARM loop.
    • Since yesterday, we have been lowering the PRCL UGF using the servos to be about 70Hz, to give us more gain margin at the high end of the phase bubble, but we still see this ringing. 
  • Two or three times my arm power buildup at 0 CARM offset, 25% MICH offset was at 20 or 21.  Then, a few other times I was only getting about 10 (which is what we have been seeing the last few days.)
    • I'm running the ASS between each lock, although I'm not running it for a full ~2 minutes or so usually. 
    • I think that the reason I was able to get to such high arm powers was excellent alignment, so maybe it's worth sitting and waiting for the ASS to have a full 2 or 3 minutes between locks, even if the ASS error signals look zero-ed out.
    • This is still a factor of 2 lower than we expect for 25% MICH offset, but the whole factor of 5 isn't explained by some mysterious loss.  At least half of it is alignment.
  • The PRCL ASC feedforward still isn't working, but I'd like to try Q's trans qpd ASC soon, with the full lock.  I think the system is ready, but scripts are not, so Q has to be here to run it.

See first plot below for the PRCL->CARM coupling just before lockloss.  The second plot is the corresponding lockloss, where the PRCL loop is starting to see that oscillation again, and it's just barely starting to get into CARM. 

  10961   Fri Jan 30 11:37:20 2015 manasaFrogsTreasureSP table madness ends

SP table has been a mess because Q and I had let our SURF leave without cleaning up.

I cleaned up the SP table, put things back where they belong and did some sorting. I will put back the optomechanics where they belong sometime later.

For now, check out the SP table next time you are looking for a Y1  or lens or BS.

 

 

 

  10960   Fri Jan 30 03:12:15 2015 diegoUpdateLSCCARM on REFL11I

[Jenne, Diego]

Tonight we continued following the plan of last night: perform the transition of CARM to REFL11_I while on MICH offset at -25%:

  • we managed to do the transition several times, keeping the UGF servos on for MICH and PRCL but turning off the DARM and CARM ones, because their contribution was rather unimportant and we feared that their excitations could affect negatively the other loops (as loops tend to see each other's excitation lines);
  • we had to tweak the MICH and PRCL UGF servos:
    • the excitation frequency for MICH was lowered to ~41 Hz, while PRCL's one was lowered to ~50 Hz;
    • PRCL's amplitude was lowered to 75 because it was probably too high and it affected the CARM loop, while MICH's one was increased to 300 because during the reduction of the CARM offset it was sinking into the noise; after a few tries we can say they don't need to be tweaked on the fly during the procedure but can be kept fixed from the beginning;
    • after the transition to REFL11_I for CARM, we engaged also its UGF servo, still at the highest frequency of the lot (~115 Hz) and with relatively low amplitude (2), to help keeping the loop stable;
    • as DARM was still on ALS, we didn't engage its UGF servo during or after the transition, but we just hold its output from the initial part of the locking sequence (after we lowered its frequency to 100 Hz;
  • however, at CARM offset 0 our arm power was less that what we had yesterday: we managed to get higher than ~8, but after Koji tweaked the MC alignment we reached ~10; we still don't understand the reason of the big difference with respect to what the simulations show for MICH offset at 25% (arm power ~50);
  • after the CARM transition to REFL11_I we felt things were pretty stable, so we tried to reduce the MICH offset to get us in the ~ -10% range, however we never managed to get past ~ -15% before losing lock, at arm power around 20;
  • we lost lock several times, but for several different reasons (IMC lost lock a couple of times, PRCL noise increased/showed some ringing, MICH railed) but our main concern is with the PRCL loop:
    • we took several measurements of the PRCL loop: the first one seemed pretty good, and it had a bigger phase bubble than usual; however, the subsequent measurements showed some weird shapes we struggle to find a reason for; these measurements were taken at different UGF frequencies, so maybe it is worth looking for some kind of correlation; morever, in the two weird measurements the UGFs are not where they are supposed to be, even if the servo was correctly following the input (or so it seemed); the last measurement was interrupted just before we lost lock because of PRCL itself;
    • we noticed a few times during the night that the PRCL loop noise in the 300-500 Hz range increased suddenly and we saw some ringing; at least a couple of times it was PRCL who threw us out of lock; this frequency range is similar to the 'weird' range we found in our measurements, so we definitely need to keep an eye on PRCL on those frequencies;
  • in conclusion, the farthest we got tonight was CARM on REFL11_I at 0 offset, DARM at 0 offset still on ALS and MICH at ~ 15% offset, arm power ~20.

 

Attachment 1: PRCL_29Jan2015_Weird_Shape.pdf
PRCL_29Jan2015_Weird_Shape.pdf
Attachment 2: ArmPowers20_MICHoffsetBeingReduced_0CARMoffset_29Jan2015.pdf
ArmPowers20_MICHoffsetBeingReduced_0CARMoffset_29Jan2015.pdf
  10959   Fri Jan 30 02:57:03 2015 JenneUpdateModern ControlFirst try with PRCL ASC Wiener filtering

[Aside - How do you rotate plots in the new elog?  It's showing them correctly in the attachments list below the entry, but not in the body of the log :( ]

I tried a round of PRCL ASC Wiener filtering today, but something wasn't right.  I was able to either make the cavity motion worse, or completely throw the cavity out of lock.  Making it less noisy didn't happen.

I took only 9 minutes of data the other day, since the PRMI didn't want to stay locked while it was daytime.  So, this wasn't a whole lot to train on.  But, even so, I designed some Wiener filters.  The plots with the designs show the calculated Wiener filter ("Wiener") and the result from vectfit ("Fit"). Below the bode plot is the coherence between that witness (seismometer direction) and the degree of freedom (QPD channel).  The fits were weighted by the choherence, so you can see that in the areas where the coherence was good, the fit was good.  Elsewhere, it's not so great.

 

Using these filters, and assuming a Cheby1 2nd order lowpass filter at 30Hz, I predicted the following residuals:

After discovering that these filters didn't work, I went rogue and also put in a high pass filter at 0.1 Hz, but that didn't really change anything.

Here is a plot of what happened in Yaw.  The Wiener filters' gains were all 0.3 here, which made the cavity motion larger, but not so large that it lost lock.  The filters ought to have gains of 1 - the Wiener calculation should figure out the gains appropriately, if I've given it enough information.  Here, as in the prediction plots above, red is the reference with the Wiener off, and black is with the Wiener filters on.  Black is supposed to be below red, if the filters are working.  Blue is the estimate of the angular motion that is being fed forward to the PRM, and you can see that at least the general shape is correct.  I do need to figure out what the resonance in the blue trace is from - it's at the same frequency as a peak in the T-240's spectrum (that I didn't save).  I suspect the cable might be touching the spaghetti pot on the inside, and making a mechanical short to pot vibrations.

Anyhow, more work to be done.  I left the PRMI locked for a while this afternoon, starting at 5:15ish, so I'll see tomorrow how long of a lock stretch I was able to capture for training. 

 

Attachment 1: WienerFilter_Witness1.png
WienerFilter_Witness1.png
Attachment 3: WienerFilter_Witness3.png
WienerFilter_Witness3.png
Attachment 9: 29Jan2015_Yaw_didnotwork.pdf
29Jan2015_Yaw_didnotwork.pdf
  10958   Thu Jan 29 17:20:58 2015 manasaConfigurationIOOX Trans Table less crazy but not enough yet

[Koji, Manasa]

We cleared up some optics and optomechanics at the X end table that are not being used and moved them to the SP table. [Ed by KA: They seemed to be leftover of the other projects. I blame them]

  10957   Thu Jan 29 16:14:10 2015 manasaUpdateGeneralUpdated to do list for FOL

Since there will be some work that supposedly will involve moving stuff on the table and racks; here is the updated to-do list for FOL.

  10956   Thu Jan 29 11:59:48 2015 manasaUpdateGeneralBeat note frequency discrepancy

Below is the set of plots comparing measurements for the green and IR beat notes frequencies. The measurements were made on the spectrum analyzer at the same time. So I have not taken measurement error into account. 
From the plots, the discrepancy is not very large.

Attachment 1:
Shows the two sets of measurements scaterred along y=x.

Attachment 2:
Since plot 1 shows the points tightly scattered to y=x, I plotted the difference between the two measurements against their mean to blow out the deviations.

I will do the same comparison using the frequency counter readout once we have RF amplifiers installed.

Attachment 1: Freq_compare1.png
Freq_compare1.png
Attachment 2: Freq_compare2.png
Freq_compare2.png
  10955   Thu Jan 29 10:14:21 2015 manasaUpdateGeneralX end space issues

It is certain that we have space issues at the X end that has been preventing us from sticking in a lens to couple light into the fiber. 

The only way out is to install a platform on the table where we can mount the lens. I have attached the a photo of how things look like at the X end (attachment 1) and also a drawing of the platform that which can hold the lens (attachment 2). Additional support to the raised platform will be added depending on how much space we can clear up on the table by moving the clamping forks of the doubler.
Steve and I have been able to gather parts that can be put together into something similar to what is shown in the drawing. 

Proposed modifications to the X end table:

1. The side panels of the table enclosure will come out while putting in the new platform.

2. The clamping forks for the doubling crystal will be moved.

Let me know of any concerns about the proposed solution. 

Attachment 1: photo.png
photo.png
Attachment 2: drawing.png
drawing.png
  10954   Thu Jan 29 09:50:47 2015 manasaUpdateElectronicsIOO rack amplifier panel

The RF amplifier panel on the IOO rack (Attachment 1) will be used to also hold the RF amplfiers for the frequency counters. The amplifiers mounted on it right now are getting +15 (orange wire) and GND (black wire) from the rack power supply (Attachment 2).

Proposed plan to install RF amplifiers:

1. Disconnect the D sub connector that powers the amplifiers and pull out the panel. The rack power supplies will NOT be shut down for this. 

2. Mount the RF amplifiers with bypass capacitors. There will be 4 amplifiers ZFL-500LN mounted on the same panel (2 for each frequency counter).

3. While putting back the panel on the IOO rack, we will need to shut down the +15V and -15V sources to connect the amplifiers to the rack power supply.

I will do this over this weekend so that we dont lose any locking time. If anybody has any concerns, let me know

Attachment 1: panel_front.jpg
panel_front.jpg
Attachment 2: panel_back.jpg
panel_back.jpg
  10953   Thu Jan 29 04:27:35 2015 ericqUpdateLSCCARM on REFL11

[ericq, Diego]

Tonight, we transitioned CARM from ALS directly to REFL11 I at 25% Mich Offset. yes

We attempted the transition twice, the first time worked, but we lost lock ~5 seconds after full transition due to a sudden ~400Hz ringup (see attached lockloss plot). The second barfed halfway, I think because I forgot to remove the CARM B offset from the first time frown

The key to getting to zero CARM offset with CARM and DARM on ALS is ekeing out every bit of PRMI phase margin that you can. Neither MICH nor PRCL had their RG filters on and I tweaked the MICH LP to attenuate less and give back more phase (the HF still isn't the dominant RMS source.) PRCL had ~60 degrees phase margin at 100Hz UGF, MICH had ~50 deg at 47Hz UGF. The error signals were comparitively very noisy, but we only cared that they held on. Also important was approaching zero slooooooooowly, and having the CARM and DARM UGF servo excitations off, because they made everything go nuts. Diego stewarded the MICH and PRCL excitation amplitudes admirably. 

Oddly, and worringly, the arm powers at zero CARM offset were only around 10-12. Our previous estimations already include the high Xarm loss, so I'm not sure what's going on with this. Maybe we need to measure our recycling gain?

I hooked up the SR785 by the LSC rack to the CM board after the first success. For the second trial, I also took TFs with respect to CM slow, but they looked nowhere near as clean as the normal REFL11 I channel; I didn't really check all the connections. I will be revisiting the whole AO situation soon. 

In any case, I think we're getting close...

Attachment 1: Jan29_REFL11_lockloss.png
Jan29_REFL11_lockloss.png
  10952   Wed Jan 28 23:53:24 2015 KojiSummaryASCXarm ASS fix

X-Arm ASS was fixed.
ASS_DITHER_ON.snap was updated so that the new setting can be loaded from the ASS screen.

The input and output matrices and the servo gains were adjusted as found in the attached image.
The output matrix was adjusted by looking at the static response of the error signals when a DC offset
was applied to each actuator.

The servo was tested with misalignment of the ITM, ETM, and BS. In fact, the servo restored transmission
from 0.15 to 1.

The resulting contrast after ASSing was ~99% level. (I forgot to record the measurement but the dark fringe level of ASDC was 4~5count.)

Attachment 1: 12.png
12.png
  10951   Wed Jan 28 17:39:17 2015 KojiConfigurationIOOX Trans Table less crazy but not enough yet

The X-end IR Trans path was cleaned up.

I have been investigating the Xarm ASS issue. The Xarm ASS sensors behaved not so straight forward.
I went to the X-end table and found some suspect of clipping and large misalignmnet in the IR trans path.
Facing with the usual chaos of the end table, I decided to clean-up the IR trans path.

The optical layout is now slightly better. But the table is, in general, still dirty with bunch of stray optics,
loose cables and fibers. We need more effort to make the table maintained in a professional manner.


- Removed unnecessary snaking optical path. Now the beam from the 1064/532 separator is divided by a 50-50 BS before the QPD without
any other steering mirrors. This means the spot size on the QPD was changed as well as the alignment. The spot on the QPD was aligned
with the arm aligned with the current (=not modified) ASS. This should be the right procedure as the spot must be centered on the end mirror
with the current ASS.

- After the 50-50 BS there is an HR steering mirror for the Thorlab PD.

- A VIS rejection filter was placed before the 50-50 BS. The reflection from the filter is blocked with a razor blade dump.

Important note to everyone including Steve:
The transmission of the VIS rejection filter at 1064nm is SUPER angular sensitive.
A slight tilt causes significant reduction of 1064nm light. Be careful.

- As we don't need double VIS filter, I removed the filter on the QPD.

- X-End QPD was inspected. There seemed large (+/-10%) gain difference between the segments.
They were corrected so that the values are matched when the beam is only on one segment.
The corrections were applied at C1:SUS-ETMX_QPDx_GAIN (x=1, 2, 3, or 4).


I decided to put "-20dB" filters on C1:SUS-ETMi_QPD_SUM and C1:SUS-ETMi_TRY (i = X or Y)
in order to make their gain to be reasonable (like 0.123 instead 0.000123 which is unreadable).
Jenne's normalization script reads relative values and the current gains instead of the absolute values.
Therefore the script is not affected.

Attachment 1: IMG_1808.JPG
IMG_1808.JPG
  10950   Wed Jan 28 17:32:26 2015 KojiUpdatePSLPMC aligned

PMC aligned.

PMC Trans increased from 0.740 to 0.782

IMC Trans increased from 16200 to 17100

  10949   Wed Jan 28 14:19:02 2015 ranaUpdateLSCTransitioned DARM to AS55Q

Why AS55/(TRX + TRY) instead of just TRX? Isn't (TRX+TRY) controlled by CARM?cool

(question is secretly from Kiwamu)

  10948   Wed Jan 28 08:53:01 2015 SteveUpdatePEMSeismometer Vertex is covered

 

Quote:

I have just put the seismometers back into their nominal positions, on the concreted slabs.  The T-240 is in the vertex, and the 2 Guralps are at the end stations.

The vertex location doesn't have a spaghetti pot right now.  There is an aluminum support for cable trays that is welded to the supports under the beam tube that is in the way.  The pot looks like it will fit barely, if it were slid totally horizontally into place.  However we can't do that with the seismometer in place.  I'll chat with Steve this afternoon about our options. 

Since I don't know that we are planning on ever putting a cable tray on the inside of the beamtube, perhaps we can cut ~6 inches of this piece away. 

Aluminum support beam removed and seismometer is covered.

  10947   Wed Jan 28 03:01:24 2015 JenneUpdateLSCTransitioned DARM to AS55Q

[Jenne, Diego]

Tonight we were able to transition DARM from DC transmission signals to AS55Q/(TRX+TRY).  That's about as far as we've gotten though.

Here are the details:

  • Set the ASDC->MICH matrix element such that the MICH fringes were 0-1.  For some reason this number seems to change by ~10% or so each night.
  • Followed main carm_cm_up script, although stopped at MICH offset of 25% (mostly because I forgot to let it go to 50% - no fundamental reason)
  • So, MICH was at 25% (with a + for the gain accidentally, even though we decided yesterday that - was better), arm powers were about 1.1 or so.
  • Took transfer functions driving DARM and looking at normalized AS55Q, and driving CARM looking at normalized REFL11I.
    • There is still not a lot of coherence in the CARM->REFL11I case, so we decided to stick with DARM for starters.
    • The TF between DARM and AS55Q looked really nice!
  • Prepared the unused DARM error signal row, including an offset before the blend matrix.
  • Transitioned over to normalized AS55Q.
    • This left the DARM servo filterbank with a zero digital offset.
    • But, the error signal had an offset before it got to the DARM filter bank.
      • This offset does not have any ramping (I don't know how to do that when building a model), so it's not as nice for reducing an offset.
      • Maybe we can, after transitioning to the new signal, move the offset to the DARM servo filterbank?
  • Was reducing the DARM offset so that we were at the true AS55Q zero crossing.
  • Saw that the UGF servo lines were starting to get a bit lost in the noise, so I increased the DARM's amplitude.
    • I don't know if the UGF servo was already too far gone and increasing the SNR couldn't recover it, or if I was driving too hard and directly kicked myself out of lock.  Either way, we lost lock.

The carm_cm_up script now should get all the way to this point by itself, although occasionally the PRMI part will lose lock (but not the arms), so you have to go back to the 3nm CARM offset and relock the central part.  I have written a little "relockPRMI.sh" script that lives in ..../scripts/PRFPMI/ that will take care of this for you.  

We were able to transition DARM to AS55Q a total of 3 or so times tonight.  The first time was with the + MICH gain, and the rest of the times were with - MICH gain.  The carm_up script now asks for a sign for the MICH gain just after asking for a CARM offset sign. 

I think that we understand all of our locklosses from these states.  Twice (including the time described above) the UGF lines got lost in the noise, and the UGF servos went crazy.  Once the PRCL loop rang up, because a filter that wasn't supposed to be on was on.  This was a terrible filter that I had made a long time ago, and was never supposed to be part of the locking sequence, but it was getting turned on by a restore script, and kept eating our phase.  Anyhow, I have deleted this terrible boost filter so it won't happen again (it was called "boost test", which may give you an idea of how non-confident I was in its readiness for prime-time).  The last time of the night I must not have been quite close enough in CARM offset, so we should probably check with a TF before making this last jump.

Diego wrote a nifty burt restoring script that will clear out all the elements of the input matrix and the normalization matrix for a row that you tell it (i.e. DARM_A will clear out all the elements in the first row of those 2 matrices).  This is useful for the switches back and forth between the _A and _B signals, to make sure that everything is in order.  So, I now run those clear scripts, then put in the elements that I want for the next step.  For example, DARM initially locks with ALS using the A row.  Then, DARM transitions to the B row for DC transmission.  Then, I prepare the A row for AS55Q locking, and I don't want any elements accidentally left from the ALS signal.  It lives in ..../scripts/LSC/InputMatrix/

 

Thoughts for tomorrow:

Daytime re-commission the Xarm ASS.

Nighttime try to get back to DARM on AS55Q and push farther forward. 

 

  10946   Tue Jan 27 21:33:39 2015 KojiUpdateASCASS retuned

I checked the situation of ASS. I wanted to know how much we are away from the maximum transmittion.

Conclusion:
ASS makes the X arm shifted from the maximum transmission. This causes the contrast degraded by ~3%.
We need to fix the Xarm ASS so that it can maximize the transmission and ignor the spot centering at ITMX.


Conditioning before the measurement:

- ASDC offset was removed
- X&Y arm was aligned by ASS

With ASS:

Average transmission: 0.86
Pmax = 1045 +/- 9 cnts
Pmin = 22 +/- 4 cnts

==> Contrast = (Pmax - Pmin)/(Pmax+Pmin) = 0.960+/-0.007

After manual alignment of the X arm (ignoring spot centering):

Average transmission: 0.88
Pmax = 1103 +/- 11 cnts
Pmin = 5 +/- 1 cnts

==> Contrast = (Pmax - Pmin)/(Pmax+Pmin) = 0.991+/-0.002

  10945   Tue Jan 27 17:58:21 2015 JenneConfigurationTreasureWelcome, Donatella!

Welcome to your new abode, Donatella!

Attachment 1: IMG_1806.JPG
IMG_1806.JPG
  10944   Tue Jan 27 15:45:26 2015 diegoUpdateLSCSmall tweaks to the locking

The UGF Servo medm page has been updated to reflect the last changes, namely the return of the sum of squares and the disappearance of Test3.

 

  10943   Tue Jan 27 14:00:05 2015 JenneUpdateModern ControlPRMI locked

PRMI is locked on sideband, starting ~4 minutes ago, to collect ASC/seismic data for feedforward.

 

  10942   Tue Jan 27 04:11:21 2015 JenneUpdateLSCSmall tweaks to the locking

[Jenne, Diego, EricQ]

We did a series of small things that may have helped with the locking, although we didn't actually get anywhere closer in CARM offset.

  • Removed the demodulation phase from the UGF servos.
    • We don't care about the phase value, just the magnitude.
    • Since we are using these through transitions between error signals, as well as with different filters coming on and off, the demodulation phase isn't constant, so the UGF servo is getting the wrong answer, and was throwing us out of lock.
    • The problems that Rana and I saw last time we had the sum of the sqrt were later discovered to be attributable to losing the peak in the noise, and hitting saturation limiters in the model, so not the fault of the sum of the squares.
    • I don't actually take the square root.  As Koji pointed out, the very next thing that happens to the signal is mag2dB, which is usually 20*log10(mag).  To compensate for the fact that I'm not taking the square root, I just use 10*log10(mag).  This removes one element of non-linearity, and leaves it at about the same number of square-ings as the complex division.
    • The excitation and measurement still happen after the multiplication.
    • The UGF screens will be updated tomorrow to reflect the change.
  • Added the new error signal rows to the LSC model's DAQ list.  So, now DARM_A_ERR and DARM_B_ERR are both acquired (and the same for CARM, MICH and PRCL).  This is to allow us to look at _DQ channels with dataviewer and DTT without having to clear the testpoints all the time.
    • We were running into too many channels being recorded, so we are not keeping the SRCL A & B signals right now.  Also, the CESAR signals that are no longer in use have been removed from the DAQ list.
  • Added a comb to the ASDC and POPDC signals, to remove the 60Hz harmonics from the MICH error signal.
    • The harmonics of the 60Hz line were dominating by more than an order of magnitude the RMS for the MICH control signal.  We couldn't afford too much phase at a few tens of Hz, so we do not notch out the original 60Hz line, although it isn't as big as, say, the 180Hz line.  So, I think that the notches are for the 2nd, 3rd, 4th and 5th order harmonics of 60Hz.  This significantly improved the RMS of the MICH output signal.
  • Lowered the MICH UGF slightly, from 48Hz to 41Hz. 
    • We wanted to go lower, to maybe 30Hz-ish, but we don't have the phase margin.  The roll mode notch is in the way, so we compromised at 41Hz.  With the comb mentioned above, the MICH control signal looks much more reasonable, and we're not injecting oodles of sensing noise into the BS.
    • Together with the comb, this ensures that we are not constantly railing the MICH output limiter, which lost us lock several times.
  • Increased the POPDC analog whitening gain from 0dB to 18dB.  We will certainly saturate POPDC when the carrier is resonant, but we were hoping to improve our SNR in our MICH error signal.  It helped noticeably, although we'll have to think about the fact that if it saturates while we're trying to acquire, the AS/POP composite signal won't be any good, and we'll kick the optics.  Also, we may have to lower the gain again as the carrier comes in to resonance.  Anyhow, something to think about.  Right now the gain is left at 18dB, and the dark offsets are set to match.
  • We may have been using the wrong side of the MICH offset, according to Q's plots.  We determined tonight that we want a (-) in the MICH gain, although the offset values can stay the same.  Even though we tried this, we didn't really get any farther in CARM offset reduction.
  • We took 2 sets of CARM and DARM loops.  They are both at 50% MICH offset, although I don't remember which sign the first one had.  The second one, at arm powers of 1.4, definitely had the new negative MICH gain. 
    • The first loop was taken at 50% MICH offset (don't remember which sign), and arm powers of about 1.15.
    • The second loop was taken at 50% MICH offset with negative gain, and arm powers of about 1.4.
    • While these numbers are not so different, maybe the first one was at roughly 300pm, and the second was at roughly 200pm, the loop shapes change dramatically.
    • The phase goes flat and we lose some of the phase margin.  Also, the magnitude is starting to get wiggly at lower frequencies.
    • See the transfer functions attached below.
  • While we were sitting at about arm powers of 1.4, with the 50% MICH offset with the negative sign in the gain, we set the REFL 11 demod phase.  It was 18.9deg (I don't remember from when), but we set it to 2.9deg to minimize the PRCL actuation in the Q-phase.  Oscillator was at 443Hz, about 10 counts. 
    • The coherence between the sqrtInvTrans signal and REFL11I looked pretty good from a few Hz to a few tens of Hz. 
    • It was late, so we decided to try transitioning over to REFL11I, and failed at the first very baby step.  So.  Ooops.
    • We should look more carefully at the TF between our current CARM signal and REFL11I.
    • We should also seriously consider using a normalized RF signal.  The SNR in the transmission PDs is just fine, although POPDC isn't a perfect choice at such high MICH offsets.
Attachment 1: CARM_27Jan2015_JCD.pdf
CARM_27Jan2015_JCD.pdf
Attachment 2: DARM_27Jan2015.pdf
DARM_27Jan2015.pdf
  10941   Mon Jan 26 21:10:04 2015 JenneUpdateModern ControlWaking up the OAF

I had a look at the OAF model today. 

Somehow, the screens that we had weren't matching up with the model.  It was as if the screens were a few versions old.  Anyhow, I found the correct screens in /userapps/oaf/common/medm, and copied them into the proper place for us, /userapps/isc/c1/medm/c1oaf.  Now the screens seem all good.

I also added 2 PCIE links between the OAF and the SUS models.  I want to be able to send signals to the PRM's pitch and yaw.  I compiled and restarted both the oaf model and the sus model.

The OAF model isn't running right now (it's got the NO SYNC error), but since it's not something that we need for tonight, I'll fix it in the morning.


My thought for trying out the OAF is to look at the coherence between seismic motion and the POP DC QPD when the PRMI is locked (no arms).  I assume that the PRM is already handled in terms of angular damping (local and oplev), so the motion will be primarily from the folding mirrors.  Then, if I can feedforward the seismometer signal to the PRM to compensate for the folding mirrors' motion, I can use the DC QPD as a monitor to make sure it's working when we're PRMI-only locked, or at low recycling gain with the arms.  But, since I'm not actually using the QPD signal, this will be independent of the arm power increase, so should just keep working.

Anyhow, that's what my game plan is tomorrow for FF.  Right now the T-240 is settling out from its move today, and the auto-zero after the move.

  10940   Mon Jan 26 17:43:52 2015 ericqConfigurationIOOMC Autolocker update

The MC autolocker hasn't been so snappy recently, and has been especially fussy today. Previously, the mcup script was triggered immediately once the transmission was above a certain threshold. However, this could waste time if it was just an errant flash. Hence, I've added a 0.5 second delay and a second threshold check before mcup is triggered. 

After breaking the lock 5ish times, it does seem to come back quicker.

  10939   Mon Jan 26 13:07:28 2015 JenneUpdatePEMSeismometers back in nominal places

I have just put the seismometers back into their nominal positions, on the concreted slabs.  The T-240 is in the vertex, and the 2 Guralps are at the end stations.

The vertex location doesn't have a spaghetti pot right now.  There is an aluminum support for cable trays that is welded to the supports under the beam tube that is in the way.  The pot looks like it will fit barely, if it were slid totally horizontally into place.  However we can't do that with the seismometer in place.  I'll chat with Steve this afternoon about our options. 

Since I don't know that we are planning on ever putting a cable tray on the inside of the beamtube, perhaps we can cut ~6 inches of this piece away. 

Attachment 1: IMG_1800.JPG
IMG_1800.JPG
Attachment 2: IMG_1805.JPG
IMG_1805.JPG
  10938   Fri Jan 23 19:38:02 2015 JenneUpdateLSCcarm_cm_up supports both signs of CARM offset

A small change, but now the carm_up script supports both sides of the CARM offset.  After the arms are locked with ALS it asks for a "+" or a "-", which indicates which sign of digital CARM offset will be added.  In the past, we have been primarily using the "+" sign.
 

  10937   Fri Jan 23 18:08:10 2015 JenneUpdateCDSEPICS freezes

So, I neglected to elog this yesterday, but yesterday we had one of those EPICS freezes that only affects slow channels that come from the fast computers.  It lasted for about 5 minutes.

Right now, we're in the middle of another - it's been about 7 minutes so far. 

Why are these happening again?  I thought we had fixed whatever was the issue.

EDIT:  And, it's over after a total of about 8 minutes.

  10936   Fri Jan 23 15:46:21 2015 ericqUpdateCDSNew netgpib scripts for SR785
Quote:

Does netgpibdata/TFSR785 work at the 40m currently? 

It does appear to work here. However, I've since supplanted TFSR785 and SPSR785 with SRmeasure, which has some simpler command line options for directly downloading a manually configured measurement. I've also set up a git repository for the gpib scripts I've done at https://github.com/e-q/netgpibdata, which could be easier than grabbing the whole 40m directory. 

  10935   Fri Jan 23 14:25:03 2015 evanUpdateCDSNew netgpib scripts for SR785

Does netgpibdata/TFSR785 work at the 40m currently? I rsynced the netgpibdata directory to LHO this morning to do some measurements, but I had to modify a few lines in order to get it to call the SR785 functions properly. My version is attached.

Attachment 1: TFSR785.zip
  10934   Fri Jan 23 10:07:15 2015 SteveUpdateSUSSUS Drift ETMX vs ETMY

ETMX YAW stopped drifting Jan 8, 2015

Quote:

I made little scripts to go with the sus driftmon buttons, that will servo the alignment sliders until the susyaw and suspit values match the references on the driftmon screen.

Attachment 1: SUSdrift30d.png
SUSdrift30d.png
  10933   Fri Jan 23 02:11:40 2015 JenneUpdateLSCCARM filters modified slightly

[Jenne, Diego]

One of tonight's goals was to tweak the CARM filters, so that we could engage the lowpass filter, to avoid the detuned double cavity pole resonance disturbing the CARM loop.

I increased the Q of the zeros in the FM3 boost so that it eats fewer than the original 18 degrees of phase at 100 Hz.  We kept losing lock though, so for each lock I backed off on the Q a little bit.  In the end, the filter eats 9 degrees of phase at 100 Hz.  I also moved the lowpass from 700 Hz to 1kHz, although that doesn't change the phase at 100 Hz very much.

We modified the carm_up script re: PRMI locking a little bit.  The PRMI is not so enthusiastic about locking immediately at 25% MICH fringe, so I backed that off.  It now acquires lock at a few percent, and then ramps up the offset.  Also, the MICH FM6 bounce roll filter is now turned on after lock is acquired, effectively giving it an extra second or two of delay beyond the rest of the filters.

We were able several times to get to some MICH offset and turn on the lowpass filter, but starting to reduce the CARM offset makes us lose lock.  I think the problem is that the UGF servo demod phase is changing as we are changing offsets, filters and error signals.  We see that the I-phase is servoed successfully to 0dB, but that the Q-phase is starting to move around by 30 degrees or more.  We either need to monitor this much more closely, and add the changing demod phases to the carm_up script, or we need to go back to the sum-of-squares situation that we had last week.  Note that we saw failures with that method for a completely separate reason:  we were getting too close to the limiters, which cause the UGF servos to glitch and the outputs jump by a significant amount.  So, the issues that we were seeing last week when we had the sum-of-squares were a different thing, that we noticed and understood later.

Anyhow, nothing too exciting and glorious tonight, but progress has been made.

Also, from some Mist simulations that Q did, Diego made a sweet plot that is now posted in the control room, so we can translate arm power to CARM offset, at various MICH offsets. 

We also took some CARM loop measurements with the new filters.  We have a little more phase than we used to, which is nice.  These traces don't have the lowpass engaged, since I was trying to see how far we could get without it.  We lost lock right after the second measurement, but I think that was to do with the UGF servos.

Attachment 2: CARM_22Jan2015.pdf
CARM_22Jan2015.pdf
  10932   Thu Jan 22 22:15:30 2015 diegoUpdateSUSCentered OpLevs

[Diego, Jenne]

We centered the OpLevs for ITMX and ITMY.
 

  10931   Thu Jan 22 18:36:11 2015 diegoUpdateIOOMC Flashes

I've been looking into the data of Jan 06 and Jan 15 taken during daytime, as the night before we left the PRC aligned in order to allow the IFO to flash; the purpose is to find out if some flashes from the IFO could propagate back to the IMC and cause it to lose lock; I will show here a sample plot, all of the others are in the archive attached.

My impression is that these locklosses of the IMC are not caused by these flashes: the signals MC_[F/L] seem quite stable until the lock loss, and I don't see any correlation with what happens to REFLDC that could cause the lockloss (apart from its drop as a consequence of the lockloss itself); in addition, in most occasions I noticed that the FSS started to go crazy just before the lock loss, and that suggests me that the lockloss source is internal to the IMC.

I can't see anything strange happen to MC_TRANS either as long as the IMC is locked, no fluctuations or weird behaviour. I also plotted the MC_REFL_SUM channel. but it is too slow to be useful for this kind of "hunt".

Attachment 1: 1104612646_zoom_1.png
1104612646_zoom_1.png
Attachment 2: elog.tar.bz2
  10930   Thu Jan 22 15:35:41 2015 diegoUpdateSUSBounce/Roll Measurements

I measured the bounce/roll frequencies for all the optics, and updated the Mechanical Resonances wiki page accordingly.

I put the DTT templates I used in the /users/Templates/DTT_BounceRoll folder; I wrote a python script which takes the exported ASCII data from such templates and does all the rest; the only tricky part is to remember to export the channel data in the order "UL UR LL" for each optic; the ordering of the optics in a single template export is not important, as long as you remember it...

Anyhow, the script is documented and the only things that may need to be modified are:

  • lines 21, 22: the "starting points" FREQ_B and FREQ_R (to accomodate noisy or bad data, as ETMX was for the Roll part in both the measurements I took);
  • line 72: the parameters of the slepian window used to average the data: the first one is the most important and indicates how much averaging will happen; more averaging means less noise but broader and lower peaks, which shouldn't be a big issue since we care only about the peak position, not its amplitude; however, if the peak is already shallow, too much averaging will make things worse instead of better;
  • lines 110, 118: the initial guess for the fit parameters;

The script is in scripts/SUS/BR_freq_finder.py and in the SVN. I attach the plots I made with this method.

Attachment 1: BR_Jan2015.tar.bz2
  10929   Thu Jan 22 03:21:24 2015 JenneUpdateLSCLocks with large MICH offsets

[Jenne, Diego, EricQ]

Tonight we worked on the acquisition sequence (including re-re-re-commissioning the UGF servos, hopefully for the last time...) for the PRFPMI with large MICH offsets. 

The procedure is all in the carm_up script, as far as things work.

We had some locklosses, but they were mostly due to non-carefulness on my part during the transitions between error signals, or the UGF servos getting upset because the oscillator peaks had gotten lost in the noise.  The one that I show here is our very last one of the night, where we are hitting the rails for the MICH signal, which is then causing the other loops to have to do weird things to try to compensate, and they lose lock.

Here also is a StripTool shot during that lock stretch.  I was in the middle of increasing the MICH offset to 75% of the fringe.  The yellow trace (called MICH_B_MON) is ASDC/POPDC normalized so that it always goes 0-1.  I was pleased to see that perhaps REFL11I and AS55Q are turning over, although as Q will tell us in a more detailed elog tomorrow, having a large MICH offset does weird things and moves the DARM zero-point.  So, maybe we aren't actually anywhere awesome yet. 

After some MICH offset, the maximum arm power is always going to be about 50, so arm powers of 8 or 10 equates to 100 pm.  We didn't get there tonight while on IR signals.


The locking sequence is now something like this:

  • Lock carm and darm on ALS, find resonances, move to 3 counts (roughly 3nm) offset.
  • Set PRMI up to acquire on REFL33I and ASDC/POPDC at 25% MICH fringe.  (After a while, I assume perhaps because the alignment is no longer tip-top, I have been by-hand reducing the MICH offset from -700counts which is 25% to -200counts, and then immediately putting it back to -700 after the PRMI acquires.)
  • Engage all 4 UGF servos
  • Reduce the CARM offset a bit, to 1.0 count, which gives arm powers of about 0.4 (with 50 being the max possible)
  • Transition CARM from ALS to sqrtInvTrans
  • Transition DARM from ALS to DC trans:  (TRY-TRX)/(TRX+TRY)
  • Reduce the oscillator amplitudes of the UGF servos
  • Reduce the CARM offset to powers of about 1
  • Ramp to 50% MICH fringe

After this, we tried a few times to lower the CARM offset, but kept losing lock, I think because the UGF servos went crazy.  The final lock, shown above, we lost because the MICH output was hitting the rails.

The problem with the MICH servo right now is the low SNR of the POPDC being used to normalize ASDC.  The control output is enormous, even if we have the 400Hz lowpass on.  We need to rethink our MICH servo, starting with a lower UGF, so that we're not injecting all this sensing noise all over the place. 


For tomorrow:

  • Re-look at MICH loop, to prevent sensing noise injection.
  • How does the large MICH offset affect our zero points for CARM and DARM?  Can we stay on DC transmission signals through 30 or 100 pm?
  • What to do next?  One or two of the locklosses were because the CARM detuned double cavity pole wasn't de-Q-ed enough, so still hit 0dB and created an unstable unity gain point.  Can we go to higher MICH offset, maybe 75%? 
  • Still need to figure out where our missing phase is for our LSC loops.  CARM and DARM are short on phase, and we could definitely use some more.  So, I will work on trying to give us filters that don't eat too much phase, but we still need to find that missing ~14 deg. 
  10928   Thu Jan 22 02:56:07 2015 ericqUpdateIOOFSS input offset adjusted

Since Rana's overhaul of the IMC, the FSS input offset had been sitting at zero. 

Over the last day or so, I had noticed the MC refl wall striptool trace looking noisier, and earlier this evening, we were suffering from a fair amount of downtime due to IMC unlocks, and failure to autolock for times on the order of ten minutes. 

While we had used ezcaservo for this in the past, I set the FSS offset manually tonight. Namely, I popped open a dataviewer trace of MC_F, and scanned the FSS offset to make MC_F go to zero. It required a good amount of offset, 4.66 V according to the FSS screen. I did this while the FSS slow servo was on, which held the FSS Fast output at zero. 

That was four hours ago; MC_F is still centered on zero, and we have not had a single IMC unlock since then. The MC refl trace is thinner too, though this may be from nighttime seismic. 

  10927   Wed Jan 21 15:27:47 2015 JenneUpdateLSCFixed LSC model bug

This problem with the CARM loop last night was the fault of a bug that I had put into the LSC model last week.  When I gave the input matrix and normalization matrix double rows, I had put the goto tags for the CARM normalization matrix rows backwards.  So, even though I thought I was not normalizing CARM, in fact I was normalizing by POPDC, which was near zero since the PRM was misaligned.

Anyhow, found, fixed, currently locking, and all seems well.

ELOG V3.1.3-